OpenStudy (anonymous):

compute the double integral y^2 dA where D is the region in the first quadrant which is enclosed by the curve defined by the equation r^2 = sin(2θ).

OpenStudy (anonymous):

not sure how to set up integral ... its from \[\int\limits_{o}^{\pi/2}\]

OpenStudy (anonymous):

then not sure how to set up the limits for the second one

OpenStudy (anonymous):

Use double angle formula on the \(\sin(2\theta)\)

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Well, actually let me think for a moment.

OpenStudy (anonymous):

I suppose you could just keep it in polar coords and convert the integrand, right?

OpenStudy (anonymous):

wouldnt that make it more complicated

OpenStudy (anonymous):

being r^2= 2sinxcosx

OpenStudy (anonymous):

\[ y^2\to r^2\sin^2(\theta) \]\[ dA=dxdy\to rdrd\theta \]

OpenStudy (anonymous):

yes thats what i have but how would i go about getting the lower and upper limits for the integral of dr

OpenStudy (anonymous):

Well, the point before was, that equation is just\[ r^2=\frac{2r\sin\theta \cdot r\cos\theta }{r^2}\to x^2+y^2=\frac{2xy}{x^2+y^2} \]

OpenStudy (anonymous):

The limits would just be \([0,r(\theta)]\)

OpenStudy (anonymous):

okay you lost me lol

OpenStudy (anonymous):

The limits of \(r\) start at \(0\) and go to \(r\) in terms of \(\theta\) typically. It depends how exactly they define the region though. You should draw the region.

OpenStudy (anonymous):

i should explain where am lost am lost where u got r^2=2rsin{theta}rcos{theta}/r^2

OpenStudy (anonymous):

Multiplied top and bottom by \(r^2\). The point was to show how it would be converted to \(x,y\)

OpenStudy (anonymous):

i see the conversion

OpenStudy (anonymous):

Regardless, try to draw it.

OpenStudy (anonymous):

its a loop on first quadrant

OpenStudy (anonymous):

thaks

OpenStudy (anonymous):

Did you figure it out?